# LIME: Less Is More for MLLM Evaluation

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) are measured on numerous benchmarks like image captioning, visual question answering, and reasoning. However, these benchmarks often include overly simple or uninformative samples, making it difficult to effectively distinguish the performance of different MLLMs. Additionally, evaluating models across many benchmarks creates a significant computational burden. To address these issues, we propose LIME (Less Is More for MLLM Evaluation), a refined and efficient benchmark curated using a semi-automated pipeline. This pipeline filters out uninformative samples and eliminates answer leakage by focusing on tasks that require image-based understanding. Our experiments show that LIME reduces the number of samples by 76% and evaluation time by 77%, while more effectively distinguishing between models. Notably, we find that traditional automatic metrics like CIDEr are insufficient for evaluating MLLMs’ captioning performance, and excluding the caption task score yields a more accurate reflection of overall model performance. All code and data are available at https://github.com/kangreen0210/LIME.
@@ 1 Introduction
In order to better understand the model’s capabilities and guide to address the shortcomings of MLLMs, researchers have developed numerous benchmarks for various tasks (Antol et al., 2015; Wei et al., 2023; Fu et al., 2023; Yue et al., 2024; Wu et al., 2024a). These benchmarks thoroughly explore the capabilities of MLLMs in various tasks such as image captioning, image question answering, and Multimodal retrieving.
However, existing MLLM benchmarks and unified evaluation frameworks cannot effectively and efficiently reflect the ability of MLLMs. Current benchmarks include numerous relatively simple samples (i.e., how many chairs are in the picture) and some incorrect questions caused by annotation issues. Most MLLMs will consistently perform on these samples (i.e., all correct or all wrong). Therefore, those benchmarks cannot fully reflect the gap between different MLLMs and across various tasks. Besides, the current unified Multimodal evaluation frameworks require significant computational resources, necessitating integrating much evaluation data from various benchmarks. The selection of effective evaluation data has been largely overlooked by current researchers.
To address the aforementioned issues, we propose to use a general data process pipeline and curate a LIME, which contains 9403 samples and is refined across 10 tasks within 6 domains. We select six major tasks in the Multimodal domain and use 9 MLLMs to refine those 10 benchmarks within the corresponding domain. To eliminate bias introduced by individual models, we chose 9 models as judges and filter samples based on their performance. On the one hand, we removed samples that most models answered correctly due to the fact that they cannot distinguish the capabilities among different models. On the other hand, we use a method that combines humans and MLLMs to filter out some abnormally difficult samples. Meanwhile, we use LLMs to filter out samples that could be answered directly from the context of the question. After that, we obtain a smaller yet higher-quality unified bench (i.e., LIME).
We conducted various experiments on LIME using both MLLMs and LLMs on different input settings, such as QA + image inputs, QA input (text-only input), and the QA + image description experiment. We have made several valuable findings:
@@ 2 Method
Figure 2 shows the statistics of different subtasks within our LIME benchmark. It is worth mentioning that the proportion of easy and wrong samples exceeds 30%. Out of the 10 subtasks, 6 have proportions exceeding 50%. Notably, for the POPE dataset, 95% of the data can be classified as noisy or erroneous. This indicates that existing benchmarks are filled with a large amount of low-quality data, which does not accurately reflect the true capabilities of MLLMs.
Inspired by MMStar (Chen et al., 2024a), we utilize open-source MLLMs and LLMs as the judges for filtering, specifically, we removing the existing annotation errors. The overall pipeline consists of three main stages: (1) Using open-source models as judges, (2) A semi-automated screening process, and (3) Eliminating answer leakage. Our approach aims to improve existing benchmarks by removing inaccurate and oversimplified data.
@@
To avoid potential biases that may exist in individual MLLMs, we select ten different types of open-source models as judges. To categorize the difficulty of each sample, we analyze the performance of all judge models on each question and label the difficulty based on the number of models that answered correctly. If , the question is classified as the easy set. If , it is classified as the middle set. Conversely, if , it is classified as the hard set.
@@ 2.2 Semi-automated screening process
Easy samples do not effectively differentiate the capabilities of various models, as most models can answer them correctly. Therefore, we remove the easy samples to assess model performance better.
@@ 2.3 Eliminating Answer Leakage
Although the previous two stages have filtered out potential errors and assessed the quality of the questions, we still need to address the potential issue of ANSWER LEAKAGE. Multimodal Answer Leakage can be summarized into two main categories: 1.Text Answerable Questions: The textual information contains all the necessary details to answer the question, making the corresponding visual information redundant. 2.Seen Questions: The model has encountered a specific question during training and has memorized the question along with its corresponding ground truth. As a result, some questions can be answered by pure text-based language models (LLMs) without relying on visual information. Therefore, we conduct a text-only check using pure text LLMs. Based on LLMs’ responses, we will remove the samples that can be directly answered without using the image. After that, we proportionally sample 1200 samples from these categories based on their difficulty levels, For benchmarks with fewer than 1,200 entries, we adapt all samples.
@@ 3 LIME: A comprehensive MLLMs benchmark
In this section, we propose LIME, a comprehensive benchmark for Multimodal Large Language Models (MLLMs). LIME streamlines existing mainstream benchmarks. Table 1 shows the main datasets included in our benchmark, as well as the data scale after careful pruning. For each sub-dataset, we aim to keep the size around 1k samples.
@@ 3.1 Task definition
LIME is composed of 10 open-source multimodal evaluation benchmarks, with scales ranging from 1,000 to 9,000. After our three-stage data curation, the data scale of each benchmark is significantly reduced. Figure 1 shows the number of samples removed at each stage compared to the original dataset. The amount of data removed varies at each stage, with the most being removed in the first stage, reflecting a large number of low-difficulty or data-leakage samples in the existing 9 MLLMs. Comparing the data volumes before and after the second stage of semi-automated screening, we can see that many datasets, such as OK-VQA and TextVQA, have a high rate of low-quality data leading to MLLMs’ incorrect answers. Additionally, some datasets, such as ScienceQA and AI2D, have a significant amount of data removed after the third stage, indicating that many questions in these datasets can be answered correctly by MLLMs without relying on images. The statistics of the curated data are shown in 1. LIME includes all data after the third stage of screening, and to further streamline the dataset, we have selected a subset from each screened benchmark based on difficulty, with each subset maintaining a data volume of around 1k.
@@ 4.1 Experiment Setting
To evaluate the quality of LIME, we conducted a series of experiments across various open-source and closed-source models. These experiments primarily encompass the following four settings:
Main experiment: To demonstrate the performance of LIME, mainstream open-source models are evaluated using a standardized process to reflect their overall performance differences.
Text only set: To prevent potential data leakage issues, we conducted validation experiments using text-only QA pairs. This was to verify whether LLMs could correctly answer questions based on text-only information.
Text-only question with Image Description set: Image Description (ID) refers to simple descriptions of images that represent superficial information contained within them. For most MLLMs, questions containing only superficial information are easy to answer; inversely, questions requiring complex visual inference are significantly more challenging. To further validate whether LIME can reflect the capabilities of MLLMs, we input text-only QA pairs combined with ID into LLMs and test their ability.
@@ 4.2 Baselines
We select LLaVA-1.5 (Liu et al., 2023a; b), LLaVA-1.6 (Liu et al., 2024), Tinny-LLaVA (Zhou et al., 2024), MiniCPM (Hu et al., 2024), Idefics-2 111https://huggingface.co/blog/idefics2, Deepseek-VL(Lu et al., 2024), CogVLM (Wang et al., 2023; Hong et al., 2023), XComposer-4KHD (Zhang et al., 2023), Mantis (Jiang et al., 2024), InternVL-1.5 and InternVL-2 (Chen et al., 2023; 2024b) as our MLLMs baseline, and LLaMA3, Yi, Yi-1.5 (AI et al., 2024), Qwen-1.5 (Bai et al., 2023a) and Qwen2 (Yang et al., 2024) as LLMs baseline. To ensure fairness in the evaluations, we use the unified evaluation framework provided by MLLMs-eval to conduct evaluation experiments on LIME. For models not supported by this framework, we refine the inference code provided by the model developers to make it compatible with the new models for the sake of aligning the results of different models.
@@
For most tasks included in LIME, we reference the metrics computation methods used in MLLMs-eval. Specifically, for tasks such as AI2D, ScienceQA, OCRBench, and POPE, we calculate the accuracy of the extracted responses. For tasks such as OK-VQA and TextVQA, the metric scores are calculated based on the overlap between the response and the candidate answers. For tasks like TextCaps and COCO-Caption2017, we use the CIDEr as the score. The ANLS metric is used for the infoVQA task, and the Relaxed Overall metric is used for the ChartQA task.
We calculate the sub-scores for each task category by taking a weighted average of the subtask scores, and then compute the overall score by weighted averaging the scores of all tasks except for the caption tasks. The details of the metrics calculation can be found in the appendix.
@@ 5.1 Main Result
| InternVL-2 2023 | 40B | 64.15 ( - ) | 80.31 | 51.69 | 48.72 | 81.12 | 75.92 | 34.90 | 56.02 |
| Qwen2-VL 2023b | 7B | 62.53 ( 1) | 79.14 | 53.05 | 51.37 | 80.83 | 62.08 | 35.70 | 89.67 |
| InternVL-1.5 2024b | 26B | 61.97 ( 1) | 79.49 | 51.69 | 52.68 | 78.96 | 63.32 | 27.90 | 90.93 |
| InternVL-2 2023 | 26B | 61.46 ( - ) | 78.82 | 54.63 | 45.64 | 79.12 | 70.54 | 32.70 | 66.54 |
| LLaVA-OneVision 2024 | 7B | 60.25 ( - ) | 78.71 | 52.37 | 51.27 | 74.50 | 66.77 | 22.00 | 106.46 |
| InternVL-2 2023 | 8B | 59.49 ( - ) | 77.84 | 49.21 | 45.15 | 76.00 | 68.54 | 32.50 | 34.00 |
| XComposer2-4KHD 2023 | 7B | 55.64 ( 4) | 71.93 | 46.28 | 44.22 | 73.29 | 58.38 | 24.40 | 87.57 |
| InternVL-2 2023 | 4B | 54.84 ( 1) | 73.97 | 47.18 | 39.89 | 71.21 | 63.31 | 30.90 | 28.83 |
| CogVLM-2 2024 | 19B | 52.02 ( 6) | 69.93 | 51.02 | 37.19 | 69.92 | 54.00 | 31.40 | 28.84 |
| Qwen2-VL 2023b | 2B | 51.59 ( 5) | 70.86 | 50.79 | 43.78 | 66.25 | 46.38 | 31.30 | 88.39 |
| InternVL-2 2023 | 2B | 51.25 ( 2) | 73.00 | 50.79 | 40.71 | 62.88 | 56.54 | 31.00 | 47.27 |
| CogVLM-1 2023 | 17B | 49.55 ( 1) | 71.34 | 55.10 | 51.45 | 59.46 | 36.54 | 19.30 | 33.92 |
| Cambrian 2024 | 34B | 48.79 ( 5) | 73.26 | 49.44 | 39.66 | 57.50 | 60.23 | 18.00 | 4.62 |
| Cambrian 2024 | 13B | 47.07 ( 4) | 72.39 | 50.79 | 41.53 | 56.04 | 49.23 | 19.50 | 6.96 |
| Cambrian 2024 | 8B | 46.42 ( 3) | 71.84 | 49.89 | 42.12 | 53.55 | 49.46 | 19.80 | 6.13 |
| InternVL-2 2023 | 1B | 45.90 ( 2) | 68.46 | 52.82 | 36.46 | 56.04 | 47.92 | 29.90 | 14.19 |
| LLaVA-1.6 2024 | 34B | 42.74 ( 3) | 67.22 | 47.00 | 30.80 | 53.21 | 53.08 | 17.10 | 66.25 |
| MiniCPM-LLaMA3-2.5 2024 | 8B | 40.63 ( 3) | 71.22 | 43.10 | 43.55 | 58.55 | 6.60 | 25.70 | 35.89 |
| LLaVA-OneVision 2024 | 0.5B | 39.88 ( 4) | 65.65 | 48.98 | 35.87 | 48.04 | 36.23 | 19.70 | 93.34 |
| LLaVA-LLaMA3 2023 | 8B | 39.82 ( 3) | 69.74 | 44.24 | 37.36 | 43.33 | 45.56 | 13.90 | 74.03 |
| Mantis-Idefics-2 2024 | 8B | 38.11 ( - ) | 66.91 | 44.24 | 36.79 | 39.75 | 43.69 | 14.80 | 82.44 |
| Deepseek-VL 2024 | 7B | 37.20 ( 2) | 65.62 | 48.50 | 34.90 | 38.50 | 44.23 | 11.70 | 68.72 |
| LLaVA-1.6-vicuna 2024 | 13B | 35.95 ( 4) | 67.29 | 43.10 | 30.00 | 41.63 | 41.54 | 14.70 | 62.23 |
| Idefics-2 2024 | 8B | 34.88 ( 2) | 66.73 | 42.00 | 46.05 | 18.50 | 47.46 | 19.60 | 77.87 |
| LLaVA-1.6-vicuna 2024 | 7B | 29.30 ( - ) | 64.80 | 41.10 | 25.75 | 32.88 | 31.77 | 10.90 | 62.20 |
| Mantis-SigLIP 2024 | 8B | 28.76 ( 1) | 58.96 | 45.60 | 29.39 | 25.79 | 35.77 | 4.90 | 74.69 |
| MiniCPM 2024 | 1.0 | 25.63 ( 2) | 56.18 | 44.00 | 21.60 | 24.58 | 35.46 | 6.70 | 72.80 |
| LLaVA-1.5 2023a | 13B | 20.17 ( 2) | 59.58 | 36.60 | 25.80 | 8.96 | 31.08 | 2.70 | 74.81 |
| LLaVA-1.5 2023a | 7B | 17.03 ( 1) | 57.27 | 32.51 | 19.97 | 7.17 | 29.81 | 2.20 | 72.47 |
| InstructBLIP-vicuna 2023 | 7B | 15.39 ( - ) | 47.87 | 45.10 | 16.75 | 6.04 | 24.77 | 2.00 | 77.61 |
| Tiny-LLaVA-1 2024 | 1.4B | 13.83 ( - ) | 34.30 | 37.00 | 9.80 | 8.33 | 27.85 | 1.60 | 61.05 |
@@
As shown in Tab 2, we evaluate the mainstream MLLMs on our LIME benchmark. Overall, models with larger parameter sizes and newer versions tend to have higher overall scores. The InternVL-1.5, InternVL-2-Large (26B,40B) and LLaVA-OneVision-7B achieve the best overall performance, whose overall scores all surpass 60%. The performance of InternVL-2-Small(1B-8B), CogVLM series, and Cambrian series are second only to InternVL-2-Large and LLaVA-OneVision-7B, with their overall scores ranging from 45% to 60%. Comparing the overall score of LIME and Origin benchmarks, it can be observed that certain model series, such as Cambrian and LLaVA-1.5, their overall scores experience a decline. Conversely, the CogVLM and LLaVA-OneVision series’ overall scores show an improvement. Specifically, that of CogVLM2 and XComposer-4KHD experience significant fluctuations, improved by 4% and 6% respectively
@@
As for the OCR task, all those MLLMs performance are all below 35%, and the performance of most models falls between 10% and 30%. This indicates that OCR tasks remain challenging for current MLLMs. Regarding with POPE subtask and caption subtask, most models demonstrate good performance. These tasks involve generating or assessing descriptions of the content in images, which proves that current MLLMs have strong image content recognition capabilities. As for the VQA task, current MLLMs have a relative high performance on the TextVQA, ChatQA and ScienceQA, which design questions directly ask the facts in the picture. While their performances are relative less on OK-VQA, infoVQA and AI2D, which require to use extra commonsense knowledge or complex reasoning to answer question. This demonstrates that the current MLLMs exhibit significant image content recognition capabilities. However, their ability to perform complex reasoning by utilizing additional knowledge is not strong. We believe this limitation may be due to the constraints of the language model component of MLLMs.

### v3

# casestudyfigurescsfList of Case Study Figures LIME: Less Is More for MLLM Evaluation

@@ Abstract
Multimodal Large Language Models (MLLMs) are measured on numerous benchmarks like image captioning, visual question answer, and reasoning. However, these benchmarks often include overly simple or uninformative samples, making it difficult to effectively distinguish the performance of different MLLMs. Additionally, evaluating models across many benchmarks creates a significant computational burden. To address these issues, we propose LIME (Less Is More for MLLM Evaluation), a refined and efficient benchmark curated using a semi-automated pipeline. This pipeline filters out uninformative samples and eliminates answer leakage by focusing on tasks that require image-based understanding. Our experiments show that LIME reduces the number of samples by 76% and evaluation time by 77%, while it can more effectively distinguish different models’ abilities. Notably, we find that traditional automatic metrics like CIDEr are insufficient for evaluating MLLMs’ captioning performance, and excluding the caption task score yields a more accurate reflection of overall model performance. All code and data are available at https://github.com/kangreen0210/LIME.
@@ 1 Introduction
In order to better understand the model’s capabilities and guide addressing the shortcomings of MLLMs, researchers develop numerous benchmarks for various tasks (Antol et al., 2015; Wei et al., 2023; Fu et al., 2023; Yue et al., 2024; Wu et al., 2024a). These benchmarks thoroughly explore the capabilities of MLLMs in various tasks such as image captioning, image question answering, and multimodal retrieving.
However, existing MLLM benchmarks and unified evaluation frameworks cannot effectively and efficiently reflect the ability of MLLMs. Current benchmarks include numerous relatively simple samples (i.e., how many chairs are in the picture) and some incorrect questions caused by annotation issues. Most MLLMs consistently perform on these samples (i.e., all correct or all wrong). Therefore, those benchmarks cannot fully reflect the gap between different MLLMs and across various tasks. Besides, the current unified multimodal evaluation frameworks require significant computational resources, necessitating integrating much evaluation data from various benchmarks. The selection of effective evaluation data is largely overlooked by current researchers.
As shown in Figure 1, to address the aforementioned issues, we propose to use a general data process pipeline and curate a LIME, which contains 9403 samples and is refined across 10 tasks within 6 domains. We select six major tasks in the multimodal domain and use 9 MLLMs to refine those 10 benchmarks within the corresponding domain. To eliminate bias introduced by individual models, we choose 9 models as judges and filter samples based on their performance. On the one hand, we remove samples that most models answer correctly due to the fact that they cannot distinguish the capabilities among different models. On the other hand, we use a method that combines humans and MLLMs to filter out some abnormally difficult samples. Meanwhile, we use LLMs to filter out samples that can be answered directly from the context of the question. After that, we obtain a smaller yet higher-quality unified bench (i.e., LIME). We conduct various experiments on LIME using both MLLMs and LLMs on different input settings, such as QA + image inputs, QA input (text-only input), and the QA + image description experiment. We make several valuable findings:
@@ 2 Method
Figure 2 shows the statistics of different subtasks within our LIME benchmark. It is worth mentioning that the proportion of easy and wrong samples exceeds 30Out of the 10 subtasks, 6 have proportions exceeding 50%. Notably, for the POPE dataset, 95% of the data can be classified as noisy or erroneous. This indicates that existing benchmarks are filled with a large amount of low-quality data, which does not accurately reflect the true capabilities of MLLMs.
Inspired by MMStar (Chen et al., 2024a), we utilize open-source MLLMs and LLMs as the judges for filtering, specifically, we remove the existing annotation errors. The overall pipeline consists of three main stages: (1) Using open-source models as judges, (2) A semi-automated screening process, and (3) Eliminating answer leakage. Our approach aims to improve existing benchmarks by removing inaccurate and oversimplified data.
@@
To avoid potential biases that may exist in individual MLLMs, we select ten different types of open-source models as judges. To categorize the difficulty of each sample, we analyze the performance of all judge models on each question and label the difficulty based on the number of models that answer correctly. We define as the number of models that correctly answer the sample. If , the question is classified as the easy set. If , it is classified as the middle set. Conversely, if , it is classified as the hard set.
@@ 2.2 Semi-automated screening process
Easy samples do not effectively differentiate the capabilities of various models, as most models can answer them correctly. Therefore, we remove the easy samples to better assess model performance.
@@ 2.3 Eliminating Answer Leakage
Although the previous two stages have filtered out potential errors and assessed the quality of the questions, we still need to address the potential issue of ANSWER LEAKAGE. Multimodal Answer Leakage can be summarized into two main categories: 1.Text Answerable Questions: The textual information contains all the necessary details to answer the question, making the corresponding visual information redundant. 2.Seen Questions: The model has encountered a specific question during training and has memorized the question along with its corresponding ground truth. As a result, some questions can be answered by pure text-based language models (LLMs) without relying on visual information. Therefore, we conduct a text-only check using pure text LLMs. Based on LLMs’ responses, we remove the samples that can be directly answered without using the image. After that, we proportionally sample 1200 samples from these categories based on their difficulty levels. For benchmarks with fewer than 1,200 entries, we adapt all samples.
@@ 3 LIME: A comprehensive MLLMs benchmark
In this section, we propose LIME, a comprehensive benchmark for Multimodal Large Language Models (MLLMs). LIME streamlines existing mainstream benchmarks. Tab 1 shows the main datasets included in our benchmark, as well as the data scale after careful pruning. For each sub-dataset, we aim to keep the size around 1k samples.
@@ 3.1 Task definition
LIME is composed of 10 open-source multimodal evaluation benchmarks, with scales ranging from 1,000 to 9,000. After our three-stage data curation, the data scale of each benchmark is significantly reduced. Figure 1 shows the number of samples removed at each stage compared to the original dataset. The amount of data removed varies at each stage, with the most being removed in the first stage, reflecting a large number of low-difficulty or data-leakage samples in the existing 9 MLLMs. Comparing the data volumes before and after the second stage of semi-automated screening, we can see that many datasets, such as OK-VQA and TextVQA, have a high rate of low-quality data leading to MLLMs’ incorrect answers. Additionally, some datasets, such as ScienceQA and AI2D, have a significant amount of data removed after the third stage, indicating that many questions in these datasets may contain potential answer leakage. The statistics of the curated data are shown in Tab 1.
@@ 4.1 Experiment Setting
To evaluate the quality of LIME, we conduct a series of experiments across various open-source and closed-source models. These experiments primarily encompass the following three settings:
Main experiment: To demonstrate the performance of LIME, we evaluate mainstream open-source and closed-source models using a standardized process to reflect their overall performance differences.
Text-only set: To prevent potential data leakage issues, we conduct validation experiments using text-only QA pairs. This verifies whether LLMs can correctly answer questions based on text-only information.
Text-only question with Image Description set: Image Description (ID) refers to simple descriptions of images that represent superficial information contained within them. For most MLLMs, questions containing only superficial information are easy to answer; however, questions requiring complex visual inference are significantly more challenging. To further validate whether LIME can reflect the capabilities of MLLMs, we input text-only QA pairs combined with ID into LLMs and test their ability.
@@ 4.2 Baselines
We select LLaVA-1.5 (Liu et al., 2023a; b), LLaVA-1.6 (Liu et al., 2024), Tinny-LLaVA (Zhou et al., 2024), MiniCPM (Hu et al., 2024), Idefics-2 111https://huggingface.co/blog/idefics2, Deepseek-VL(Lu et al., 2024), CogVLM (Wang et al., 2023; Hong et al., 2023), XComposer-4KHD (Zhang et al., 2023), Mantis (Jiang et al., 2024), InternVL-1.5 and InternVL-2 (Chen et al., 2023; 2024b) as our MLLMs baseline, and LLaMA3, Yi, Yi-1.5 (AI et al., 2024), Qwen-1.5 (Bai et al., 2023a) and Qwen2 (Yang et al., 2024) as LLMs baseline. To ensure fairness in the evaluations, we use the unified evaluation framework provided by lmms-eval (Zhang et al., 2024b) to conduct evaluation experiments on LIME. For models not supported by lmms-eval, we refine the inference code provided by the model developers to make it compatible with the new models for the sake of aligning the results of different models.
@@
For most tasks included in LIME, we reference the metrics computation methods used in lmms-eval. Specifically, for tasks such as AI2D, ScienceQA, OCRBench, and POPE, we calculate the accuracy of the extracted responses. For tasks such as OK-VQA and TextVQA, we calculate the metric scores based on the overlap between the response and the candidate answers. For tasks like TextCaps and COCO-Caption2017, we use CIDEr as the score. The ANLS metric is used for the infoVQA task, and the Relaxed Overall metric is used for the ChartQA task.
We calculate the sub-scores for each task category by taking a weighted average of the subtask scores, and then compute the overall score by weighted averaging the scores of all tasks except for the caption tasks. The details of the metrics calculation are provided in Tab 7.
@@ 5.1 Main Result
| GPT-4O | - | 52.63 | - | 47.18 | 42.95 | 57.63 | 56.15 | 72.39 | 47.84 |
| claude-3-5-sonnet | - | 51.99 | - | 35.89 | 50.33 | 56.38 | 44.69 | 73.91 | 28.00 |
| Gemini-1.5-Pro-Vision | - | 49.46 | - | 54.63 | 37.71 | 55.33 | 50.15 | 73.26 | 41.38 |
| GPT-4-Vision-Preview | - | 42.23 | - | 42.44 | 33.86 | 48.00 | 42.39 | 55.22 | 29.14 |
| InternVL-2 2023 | 40B | 66.85 ( - ) | 80.31 | 51.69 | 48.72 | 81.12 | 75.92 | 75.87 | 56.02 |
| Qwen2-VL 2023b | 7B | 65.28 ( 1) | 79.14 | 53.05 | 51.37 | 80.83 | 62.08 | 77.61 | 89.67 |
| InternVL-1.5 2024b | 26B | 64.12 ( 1) | 79.49 | 51.69 | 52.68 | 78.96 | 63.32 | 60.65 | 90.93 |
| InternVL-2 2023 | 26B | 63.98 ( - ) | 78.82 | 54.63 | 45.64 | 79.12 | 70.54 | 71.09 | 66.54 |
| InternVL-2 2023 | 8B | 62.00 ( 1 ) | 77.84 | 49.21 | 45.15 | 76.00 | 68.54 | 70.65 | 34.00 |
| LLaVA-OneVision 2024 | 7B | 61.95 ( 1 ) | 78.71 | 52.37 | 51.27 | 74.50 | 66.77 | 47.83 | 106.46 |
| XComposer2-4KHD 2023 | 7B | 57.52 ( 4) | 71.93 | 46.28 | 44.22 | 73.29 | 58.38 | 53.04 | 87.57 |
| InternVL-2 2023 | 4B | 57.22 ( 1) | 73.97 | 47.18 | 39.89 | 71.21 | 63.31 | 67.17 | 28.83 |
| CogVLM-2 2024 | 19B | 54.44 ( 6) | 69.93 | 51.02 | 37.19 | 69.92 | 54.00 | 68.26 | 28.84 |
| Qwen2-VL 2023b | 2B | 54.00 ( 5) | 70.86 | 50.79 | 43.78 | 66.25 | 46.38 | 68.04 | 88.39 |
| InternVL-2 2023 | 2B | 53.64 ( 2) | 73.00 | 50.79 | 40.71 | 62.88 | 56.54 | 67.39 | 47.27 |
| CogVLM-1 2023 | 17B | 51.03 ( 1) | 71.34 | 55.10 | 51.45 | 59.46 | 36.54 | 41.96 | 33.92 |
| Cambrian 2024 | 34B | 50.17 ( 5) | 73.26 | 49.44 | 39.66 | 57.50 | 60.23 | 39.13 | 4.62 |
| Cambrian 2024 | 13B | 48.57 ( 4) | 72.39 | 50.79 | 41.53 | 56.04 | 49.23 | 42.39 | 6.96 |
| InternVL-2 2023 | 1B | 48.21 ( 3) | 68.46 | 52.82 | 36.46 | 56.04 | 47.92 | 65.00 | 14.19 |
| Cambrian 2024 | 8B | 47.95 ( 4) | 71.84 | 49.89 | 42.12 | 53.55 | 49.46 | 43.04 | 6.13 |
| LLaVA-1.6 2024 | 34B | 44.06 ( 3) | 67.22 | 47.00 | 30.80 | 53.21 | 53.08 | 37.17 | 66.25 |
| MiniCPM-LLaMA3-2.5 2024 | 8B | 42.61 ( 3) | 71.22 | 43.10 | 43.55 | 58.55 | 6.60 | 55.87 | 35.89 |
| LLaVA-OneVision 2024 | 0.5B | 41.40 ( 4) | 65.65 | 48.98 | 35.87 | 48.04 | 36.23 | 42.83 | 93.34 |
| LLaVA-LLaMA3 2023 | 8B | 40.90 ( 3) | 69.74 | 44.24 | 37.36 | 43.33 | 45.56 | 30.22 | 74.03 |
| Mantis-Idefics-2 2024 | 8B | 39.25 ( - ) | 66.91 | 44.24 | 36.79 | 39.75 | 43.69 | 32.17 | 82.44 |
| Deepseek-VL 2024 | 7B | 38.10 ( 2) | 65.62 | 48.50 | 34.90 | 38.50 | 44.23 | 25.43 | 68.72 |
| LLaVA-1.6-vicuna 2024 | 13B | 37.08 ( 4) | 67.29 | 43.10 | 30.00 | 41.63 | 41.54 | 31.96 | 62.23 |
| Idefics-2 2024 | 8B | 36.39 ( 2) | 66.73 | 42.00 | 46.05 | 18.50 | 47.46 | 42.61 | 77.87 |
| LLaVA-1.6-vicuna 2024 | 7B | 30.15 ( - ) | 64.80 | 41.10 | 25.75 | 32.88 | 31.77 | 23.70 | 62.20 |
| Mantis-SigLIP 2024 | 8B | 29.13 ( 1) | 58.96 | 45.60 | 29.39 | 25.79 | 35.77 | 10.65 | 74.69 |
| MiniCPM 2024 | 1.0 | 26.15 ( 2) | 56.18 | 44.00 | 21.60 | 24.58 | 35.46 | 14.57 | 72.80 |
| LLaVA-1.5 2023a | 13B | 20.38 ( 2) | 59.58 | 36.60 | 25.80 | 8.96 | 31.08 | 5.87 | 74.81 |
| LLaVA-1.5 2023a | 7B | 17.20 ( 1) | 57.27 | 32.51 | 19.97 | 7.17 | 29.81 | 4.78 | 72.47 |
| InstructBLIP-vicuna 2023 | 7B | 15.55 ( - ) | 47.87 | 45.10 | 16.75 | 6.04 | 24.77 | 4.35 | 77.61 |
| Tiny-LLaVA-1 2024 | 1.4B | 13.95 ( - ) | 34.30 | 37.00 | 9.80 | 8.33 | 27.85 | 3.48 | 61.05 |
@@
As shown in Tab 2, we evaluate both open-source and closed-source MLLMs using our LIME benchmark. Overall, for closed-source models, GPT-4O achieves the best performance with a score of 52%, while for open-source models, models with larger parameter sizes and newer versions tend to have higher overall scores. InternVL-1.5, InternVL-2-Large (26B, 40B), and LLaVA-OneVision-7B achieve the best overall performance, with their overall scores all surpassing 60%. The performance of InternVL-2-Small (1B-8B), the CogVLM series, and the Cambrian series follows, with their overall scores ranging from 45% to 60%.
Comparing the overall scores of LIME and Origin benchmarks, we observe that certain model series, such as Cambrian and LLaVA-1.5, experience a decline in overall scores. Conversely, the CogVLM and LLaVA-OneVision series show an improvement, with CogVLM2 and XComposer-4KHD experiencing significant increases of 4% and 6%, respectively.
@@
Tab 6 provides more detailed experimental results. Regarding caption subtasks, most models demonstrate good performance. These tasks involve generating or assessing descriptions of the content in images, which indicates that current MLLMs possess strong image content recognition capabilities. As for the VQA task, current MLLMs perform relatively well on TextVQA, ChatQA, and ScienceQA, where the questions directly ask about facts in the picture. However, their performance is relatively lower on OK-VQA, infoVQA, and AI2D, which require additional commonsense knowledge or complex reasoning to answer the questions. This demonstrates that current MLLMs exhibit significant image content recognition capabilities but are limited in their ability to perform complex reasoning using additional knowledge. We believe this limitation may be due to constraints in the language model component of MLLMs.